OpenStudy (anonymous):

Suppose Q = x^{2} y^{3} . If \frac{dQ}{dt} = -2 and \frac{dx}{dt} = -1, find \frac{dy}{dt} when x = 3 and y = 2 . \frac{dy}{dt} =

OpenStudy (anonymous):

i tried so many times and i got 1/3 but this doesnt seem to be correct

OpenStudy (experimentx):

Suppose \( Q = x^{2} y^{3}\) . If \( \frac{dQ}{dt} = -2 \) and \( \frac{dx}{dt} = -1 \), find \( \frac{dy}{dt} \) when x = 3 and y = 2 .\( \frac{dy}{dt} \) =

OpenStudy (anonymous):

4/27 ?

OpenStudy (anonymous):

@ acekly not correct

OpenStudy (experimentx):

\( \frac{dQ}{dt} = 2x y^3\;dx/dt + x^23y^2\;dy/dt \) , put the given values ....

OpenStudy (anonymous):

5/54?

OpenStudy (anonymous):

@ acekly not correct

OpenStudy (anonymous):

@experimentX Yep, I got -50/108 from that, I think, but I may have commited a mistake somewhere...my brain must be malfunctioning.

OpenStudy (anonymous):

i got 54/23

OpenStudy (anonymous):

not correct either

OpenStudy (experimentx):

@bmp don't worry i haven't slept all night

OpenStudy (anonymous):

ok its 23/54 for the answer

OpenStudy (anonymous):

i got it, thanks guys !

OpenStudy (anonymous):

23/54?

OpenStudy (anonymous):

alryte